OpenStudy (anonymous):

Solve: ap^2 +py -x =0

OpenStudy (anonymous):

ap2=x−py

OpenStudy (anonymous):

ap2+py−x=0 Move all terms not containing a to the right-hand side of the equation. ap2=−py+x Move all terms not containing a to the right-hand side of the equation. ap2=x−py

OpenStudy (anonymous):

A particle is moving along the x- axis so that its distance x, in meters, at time t in seconds, is x(t)=[3,2\sqrt{\tan5l.}\]\sqrt{\tan5t.} find its velocity at time t =1,5 seconds\]